OpenStudy (anonymous):

Evaluate m^0 - n^2 for m = 2 and n = -1

OpenStudy (anonymous):

anything to the zero power is 1. anything to the second power is itself times itself.

OpenStudy (anonymous):

= (2)^0 - (-1)^2 = 1 - 1 = 0